OpenStudy (anonymous):

If a pro basket ball player has a vertical leap of about 30 inches, what is the hang time? Use the hang time function v=48T to the second power

OpenStudy (amistre64):

since the apex of his jump is 30 inches we can use this to backtrack

OpenStudy (amistre64):

i got no idea what a hangtime function is tho

OpenStudy (anonymous):

V=48T^2 ?

OpenStudy (amistre64):

-16t^2 +v0 t = 30/12 right?

OpenStudy (anonymous):

thats the position function well close to it

OpenStudy (amistre64):

or rather the vertex of the parabola this makes is at 0,30

OpenStudy (amistre64):

.. im at a loss for it at the moment

OpenStudy (amistre64):

y = -16(t- n)^2 + 30/12 y = -16(t^2 -2t x +n^2) + 15/6 y = -16t^2 +32n t +(-16n^2 + 15/6), the ( ) part is the initial height which is 0 right? -16n^2 + 15/6 = 0 (-96n^2 + 15)/6 = 0 n^2 = 15/96 = 5/32 n = sqrt(5/32) if im right ... still aint determined if im right tho lol

OpenStudy (amistre64):

y = -16t^2 +32(5/32) t y = -16t^2 +5t = 0 t(-16t + 5) = 0 when t=0 and t=5/16 is my best guess

OpenStudy (amistre64):

its wrong of course, but then i cant make sense of the questions as stated

OpenStudy (amistre64):

a better thought might be to consider the time it takes to drop from 30 inches and double that

OpenStudy (amistre64):

-16t^2 + 30/12 = 0 -16t^2 + 15/6 = 0 -16(6)t^2/6 + 15/6 = 0 (-96t^2+15)/6 = 0 when t^2 = 15/96

OpenStudy (amistre64):

15/96 = 5/32 t = sqrt(5/32) , which should be half the time starting at the downward part of the jumpers arc thru the air, so double that to get the full time 2 sqrt(5/32)

OpenStudy (amistre64):

.7906 seconds if we are counting in seconds :) and thats my best offer